OpenStudy (anonymous):

A bag contains 10 white golf balls and 6 striped golf balls. A golfer wants to add 112 golf balls to the bag. He wants the ratio of white to striped golf balls to remain the same. How many of each should he add?

OpenStudy (jjuden):

Let w = number of added white balls Let s = number of added striped balls Final number of balls = 10 + 6 + 112 = 128 = 10 + w + 6 + s 128 = 16 + w + s w = 112 - s........................(1) The ratio of white to striped balls must remain as 10/6 Therefore we can write the following equation: 10+w6+s=106 Cross-multiplying gives: 60 + 6w = 60 + 10s Therefore 6w = 10s.............(2) Now you just need to substitute the value for w found in (1) into (2) and solve for s.

OpenStudy (jjuden):

42 striped and 70 white if u dont want to do the math